# Improving Steering and Verification in AI-Assisted Data Analysis with Interactive Task Decomposition

### v1

@@ 4.1.3. Task Decomposition: Stepwise System
Unlike the Phasewise system where each component reflects the entire task, the Stepwise system decomposes the task into subgoals, which have intermediate objectives. Each subgoal (except the first, which loads the dataset (Figure 4, a)), is represented as a pair of components: Subgoal Assumptions and Actions, and Subgoal Code and Output.
A) Subgoal Assumptions and Actions: Each subgoal starts with a short description of its objective in natural language, followed by several assumptions and actions based on the dataset or previous steps (Figure 4, b). We designed LLM prompts so that each subgoal would focus on one specific objective such as pre-processing data, filtering columns, performing calculations, and displaying plots. Users may reorder assumptions and actions, add or remove assumptions, and edit them directly. Once the user is satisfied with them, they can proceed to generate the subgoal code and output.
B) Subgoal Code and Output: Similar to Code and Output in the Phasewise system, in this component, the system generates code to solve the task based on the previous assumptions and actions (Figure 4, c). The code is immediately executed and can be edited. Once executed, the system generates the next subgoal to allow the user to either reflect on the current subgoal or start working on the next. This process continues until the task is finished and the requirements have been satisfied, in which case the next Subgoal Assumptions and Actions will indicate completion. However, if the user still wants to continue, they can add assumptions and actions to continue.
Figure 4. Overview of the Stepwise system’s task flow, which decomposes each task into subgoals containing two components. Assumptions and Actions includes the NL subgoal and editable AI assumptions b. Code and Output contains a code editor a c, a dataframe and variable inspector d, and the code output e. Section 4.1.3 details these features.
@@ 4.1.4. Editable LLM Assumptions and Actions
[…]
@@ 4.1.6. Managing Edits
Within each component, edits can be either pending or submitted. A submitted edit means that the edits have been applied to generate the next component, whereas a pending edit has not. Pending edits can be reverted using an undo button. However, once an edit is submitted, our system introduces a new branch to preserve the original, unedited version, while incorporating the edited version in the “main” branch. New branches are displayed in a tabbed ribbon at the top of the UI. To allow iteration while minimizing proliferation of branches, new branches are not created when the user edits the last generated component in the stream of components. Branching allows users to keep track of previous edits and switch between edits as needed.
Figure 5. In the Stepwise and Phasewise systems, users can select any code in the editor to ask questions a from the AI. This will create a question box to the right of the main components in which users can ask their clarification question b. The question box will then be replaced with the AI’s response c, based on the query and the selected text.
@@ 4.1.7. Side Conversations
Figure 6. The run Side Query button a is placed to the right of each code execution block in the Stepwise and Phasewise systems, and the input and output assumptions in the Phasewise system. This will open the question box b in which the user can specify a natural language data exploration query on any of the intermediary variables or the original dataset. The question box will then be replaced with the AI’s response which includes the code and any outputs, including visualizations c.
@@ 4.3. System Implementation
[…]
@@ 5.1. Participants
We recruited 18 participants (10 men, 8 women, 0 non-binary) from a large research university. Participants were pre-screened to ensure they were proficient in writing Python code, familiar with Python data science libraries, and experienced in regularly performing data analysis tasks. In terms of data analysis experience, five participants reported having 1-2 years, seven having 3-5, and six more than five years. The majority (14 participants) used Python daily, while the rest used it at least weekly. All reported familiarity with data science libraries like numpy, matplotlib, with 15 also familiar with pandas. Jupyter Notebooks were used daily by eight participants, weekly by six, monthly by two, and rarely by two. For English proficiency in technical contexts, 16 participants felt very comfortable, while two felt somewhat comfortable. In LLM usage for coding, seven reported using daily, seven weekly, and four using monthly or less.
@@ 5.2. Data Analysis Tasks
We designed six tasks derived from the ARCADE benchmark (Yin et al., 2022), which contains a diverse set of tasks from various datasets on Kaggle (Kaggle, 2023). These tasks included a series of natural language (NL) queries written by professional data scientists with the intention of interacting with an AI assistant. We selected tasks to not require specific domain knowledge, targeting for participants to solve them within a 15-minute time frame. However, we also wanted to make sure that the tasks included additional complexities that would make it difficult for the AI to correctly solve them without proper user verification and intervention. Therefore, we selected and altered tasks and their corresponding datasets with some format inconsistencies and altered distributions. For instance, Task 2 (Table 2) requires splitting tags and themes with commas before grouping tags by themes. To increase complexity, we modified the tags and themes columns to have only the first theme or tag in capital case, with the rest in lowercase. See Table 2 for details of the six study tasks. To guarantee consistent failure of the system on these tasks, we ran each query 10 times across all systems, ensuring a 100% failure rate.
@@
[…]
@@ 5.3. Study Procedure
The order of the Phasewise, Stepwise, and Conversational (baseline) systems was counterbalanced across participants and tasks to minimize order effects, while tasks were fixed from T1 to T6. Each participant began with a 10-minute introduction to their first assigned system, followed by a 10-minute warm-up task to familiarize themselves with its features and mitigate learning effects. Participants then proceeded to the main study tasks, where they were given a dataset and a NL query. The researcher explained the dataset and relevant columns for each task. Participants proceeded to execute the task and were asked to think aloud throughout the study (Fonteyn et al., 1993).
Participants were made aware that identifying and correcting mistakes made by the AI was their responsibility. They were asked to notify the experimenter once they believed they have achieved a correct result using the AI tool. The experimenter would then check their result and provide hints if necessary. Completion criteria for each task required resolving both issues listed in Table 2.
Following the completion of each task, participants were asked about their choice of method for steering the AI (e.g., editing the execution plan versus directly editing the code) and their verification processes. After completing two tasks under each condition, participants completed a questionnaire including Likert items about their ability to verify, intervene and steer the AI, sense of control, information overload, frustration levels, and the utility of specific features. Additionally, participants discussed their experience with each system in a 5-minute semi-structured interview. After completing all tasks using the three systems, a final questionnaire elicited a comparative evaluation of the systems in terms of confidence, usability, control, and ease of verification and steering.
The sessions was conducted in-person, lasting approximately 2.5 hours with a short break at the halfway point. Consent was obtained before running the study and each participant was compensated with a GBP £50 Amazon gift card. Our study protocol was reviewed and approved by our institution’s ethics and compliance review board.
@@ 5.4. Data Collection and Analysis
We recorded the audio and screen activity during each session using MS Teams. Audio recordings were transcribed for analysis. User system interactions and feature usage was also logged.
The think-aloud data was our main source of understanding how participants used the different systems and what they thought about them in comparison with each other. We transcribed the think-aloud data and post-condition interviews and two researchers performed a negotiated, directed qualitative analysis. Ahead of the analysis, we identified a focused set of research themes concerning steering and verification during the AI-assisted data analysis process, and report our findings organised by these themes in Section 6. Because we were interested in specific themes a priori, and were not developing a reusable coding scheme, our analysis differs from the more commonly applied inductive approach (Braun and Clarke, 2006). We did not develop a codebook, and this is not a situation in which it is appropriate to seek inter-rater reliability. Instead, in accordance with qualitative coding best practices, the two researchers iteratively discussed their interpretation of the findings and negotiated each disagreement until it was resolved (McDonald et al., 2019).
@@
We analyzed post-condition Likert responses to compare the three systems and determine any statistically significant differences using a Friedman Chi Square test on the responses for each question with the system type as the independent variable. When significant differences were found ($\alpha=0.05$), a Wilcoxon signed-rank test identified pairwise significant comparisons, after applying Bonferroni correction ($\alpha=0.016$).
@@ 6. Results
In this section, we present a comparative analysis of the Stepwise and Phasewise AI tools versus the Conversational baseline. Our findings are derived from study observations, log data, participant (P1-P18) think-aloud data, post-condition surveys, and post-study interviews. In turn, we present the results regarding task completion (Section 6.1), steering and control (Section 6.2), and verification (Section 6.3).
@@ 6.2.1. Steering by Directly Editing AI’s Assumptions and Actions
6.2.2. Enhanced Control Through Step-by-Step Task Decomposition
A distinct advantage observed with the Stepwise system was the enhanced control participants reported over the data analysis process. This was mainly attributed to tackling the task in smaller, manageable segments. For instance, P16 reported an increased sense of control, by being able to “easily edit the assumptions and actions in each step.” Similarly, P11 felt “much happier” and “more confident”, attributing it to the ability to “manipulate steps naturally.” P10 also shared a sense of more control over the AI, stating that they “were not scared” to make edits to what the AI was doing, as “it was just a couple of lines,” and “it was more inviting to edit the assumptions.” P17 mentioned that the Stepwise system facilitated an iterative process “where [they] could easily go back and change something”.
@@ 6.2.3. Steering by Manually Editing Code
[…]
@@ 6.2.4. Preference for Conversational Steering
In some instances, participants (P4, P5, P6, P8, and P11) specifically indicated a preference for steering the AI through natural language prompts. They favored the Conversational method of steering the AI over editing the structured assumptions, actions, or execution plans. For example, P4 expressed difficulty in understanding how changes to the assumptions in the Phasewise system, affected the LLM’s output. In contrast, P4 had a more accurate mental model of how to interact with the Conversational baseline, stating “I know exactly how writing a prompt is going to affect it.” Others felt constrained by the need to adhere to a specific structure, expressing a preference for more free-form interactions. For instance, P8 described the Conversational system as easier and faster for “directing the AI using natural language”, compared to the Stepwise system where they felt they were “trying to change the syntax of the AI.” Similarly, P5 wanted to “intentionally write vague prompts and see how much [the AI] understands.” P4 believed that the Conversational system required less cognitive effort, stating “I don’t like spending that effort to think about it.” P11 mentioned feeling “less critical” when using the Conversational tool, allowing the AI to “go and figure it out” on their behalf. P11 elaborated: “I knew what it was [that] I wanted it to consider, but when [the tool] is expecting a structured input then I was more concerned with providing it in a nice and structured manner.”
@@ 6.2.5. Avoiding Edits that Lead to Inconsistency or Regeneration
  • if upstream components would go out of sync and become invalidated because of updating downstream components,
  • updating upstream components would cause downstream components to be regenerated, ignoring any prior edits on the downstream component.
@@ 6.3.1. Verification through Reading Code and Asking Questions
[…]
@@ 6.3.2. Inspecting Intermediary Variables
All participants used the intermediary variable inspection feature available in all three systems for verification. Participants inspected variables to “compare between turns” (P1), and to “see if [the system] has done the [operation] correctly” (P14). P12 mentioned that the verification process was similar to “debugging” and P11 stated that inspecting all the variables.

### v2

@@ 4.1.3. Task Decomposition: Stepwise System
Unlike the Phasewise system where each component reflects the entire task, the Stepwise system decomposes the task into subgoals, which have intermediate objectives. Each subgoal (except the first, which loads the dataset (Figure 4, [Uncaptioned image])), is represented as a pair of components: Subgoal Assumptions and Actions, and Subgoal Code and Output.
A) Subgoal Assumptions and Actions: Each subgoal starts with a short description of its objective in natural language, followed by several assumptions and actions based on the dataset or previous steps (Figure 4, [Uncaptioned image]). We designed LLM prompts so that each subgoal would focus on one specific objective such as pre-processing data, filtering columns, performing calculations, and displaying plots. Users may reorder assumptions and actions to change their priority, add or remove assumptions, and edit them directly. Once the user is satisfied with them, they can proceed to generate the subgoal code and output.
B) Subgoal Code and Output: Similar to Code and Output in the Phasewise system, in this component, the system generates code to solve the task based on the previous assumptions and actions (Figure 4, [Uncaptioned image]). The code is immediately executed and can be edited. Once executed, the system generates the next subgoal to allow the user to either reflect on the current subgoal or start working on the next. This process continues until the task is finished and the requirements have been satisfied, in which case the next Subgoal Assumptions and Actions will indicate completion. However, if the user still wants to continue, they can add assumptions and actions to continue.
Figure 4. Overview of the Stepwise system’s task flow, which decomposes each task into subgoals containing two components. Assumptions and Actions includes the NL subgoal and editable AI assumptions Refer to caption. Code and Output contains a code editor Refer to caption Refer to caption, a dataframe and variable inspector Refer to caption, and the code output Refer to caption. Section 4.1.3 details these features.
@@ 4.1.4. Editable LLM Assumptions and Actions
We prompted the LLM to generate each assumption paired with its corresponding action in the format of <assumption> - <action>. We also prompted the LLM to enclose column names, variables, and keywords in backticks, which could be rendered into editable components highlighted with a different color. The aim of these interventions was to reduce information overload and improve the efficiency of the editing process.
@@ 4.1.5. Code Execution and Intermediary Variables
Figure 5. The tabbed ribbon displays all the branches created after editing various nodes. Users can select a different tab to switch to that branch. Each tab indicates where the edit occurred and how much it has progressed (number of total nodes).
@@ 4.1.6. Managing Edits
Within each component, edits can be either pending or submitted. A submitted edit means that the edits have been applied to either generate the next component or regenerate downstream components, whereas a pending edit has not. Pending edits can be reverted using an undo button. However, once an edit is submitted, our system introduces a new branch to preserve the original, unedited version, while incorporating the edited version in the “main” branch. New branches are displayed in a tabbed ribbon at the top of the UI as shown in Figure 5. To allow iteration while minimizing proliferation of branches, new branches are not created when the user edits the last generated component in the stream of components. Branching allows users to keep track of previous edits and switch between edits as needed.
Figure 6. In the Stepwise and Phasewise systems, users can select any code in the editor to ask questions Refer to caption from the AI. This will create a question box to the right of the main components in which users can ask their clarification question Refer to caption. The question box will then be replaced with the AI’s response Refer to caption, based on the query and the selected text.
@@ 4.1.7. Side Conversations
Figure 7. The run Side Query button Refer to caption is placed to the right of each code execution block in the Stepwise and Phasewise systems, and the input and output assumptions in the Phasewise system. This will open the question box Refer to caption in which the user can specify a natural language data exploration query on any of the intermediary variables or the original dataset. The question box will then be replaced with the AI’s response which includes the code and any outputs, including visualizations Refer to caption.
@@ 4.3. System Implementation
All three variants are built as a web application and Python server stack. The web application is written in TypeScript and the React web framework to include the interface elements described in Section 4.1. The web application interacted with the Python server for uploading datasets, obtaining their descriptive summaries, running code, and retrieving their execution results. It also called the GPT-4 Turbo models from OpenAI. Each component in the Phasewise or Stepwise system is represented as a node in a tree data structure inside the application. This enables tracing the path from each node to the root node to prepare the context prompt for interacting with the LLM and generating the next component. It also provides state management for the edits that create branches and is used to render the tabbed ribbon interface. The Monaco Editor is used as the code editor in each of the code execution blocks and for syntax highlighting the non-editable code pieces. To enable code execution on the Python server, and to retrieve code execution outputs and intermediary variables, we used the IPython kernel. We used the %matplotlib inline command which returned all plots as base64 images that could be included as a response inside the REST APIs.
@@ 5.1. Participants
We recruited 18 participants (10 men, 8 women, 0 non-binary) from a large research university. Participants were pre-screened to ensure they were proficient in writing Python code, familiar with Python data science libraries, and experienced in regularly performing data analysis tasks. In terms of data analysis experience, five participants reported having 1–2 years, seven having 3–5, and six more than five years. The majority (14 participants) used Python daily, while the rest used it at least weekly. All reported familiarity with data science libraries like numpy, matplotlib, with 15 also familiar with pandas. Jupyter Notebooks were used daily by eight participants, weekly by six, monthly by two, and rarely by two. For English proficiency in technical contexts, 16 participants felt very comfortable, while two felt somewhat comfortable. In LLM usage for coding, seven reported using daily, seven weekly, and four using monthly or less.
@@ 5.2. Data Analysis Tasks
We designed six tasks derived from the ARCADE benchmark (Yin et al., 2022), which contains a diverse set of tasks from various datasets on Kaggle (Kaggle, 2023). These tasks included a series of natural language (NL) queries written by professional data scientists with the intention of interacting with an AI assistant. We selected tasks to not require specific domain knowledge, targeting for participants to solve them within a 15-minute time frame. However, we also wanted to make sure that the tasks included additional complexities that would make it difficult for the AI to correctly solve them without proper user verification and intervention. Therefore, we selected and altered tasks and their corresponding datasets with some format inconsistencies and altered distributions. For instance, Task 2 (Table 2) requires splitting tags and themes with commas before grouping tags by themes. To increase complexity, we modified the tags and themes columns to have only the first theme or tag in capital case, with the rest in lowercase. See Table 2 for details of the six study tasks. We ran each of the six tasks (query + dataset) 10 times on all three systems, ensuring a consistent 100% failure rate.
@@
Task 1: Show me the top five highly rated products by Nivea big-basket-products.csv Not recognizing multiple sub-brands of "Nivea" (1) and not cleaning the Rating column to accurately extract numeric values (2).
Task 2: What is the most common tag associated with each theme? anime-list.csv Not correctly splitting Themes by comma (1) and overlooking the inconsistency in casing between Tags and Themes (2).
Task 3: Display the top 20 most popular drama names that have only one unique genre? Popularity is based on drama rating and votes. korean-drama.csv Not filtering genres labeled as "Unknown" (1) and extreme outliers in votes (2)
Task 4: What are the top ten positions (based on mean salary) for working remotely in US-based companies? data-science-job-salaries.csv Not cleaning Country Code (1) and not identifying remote companies using Remote Ratio (2).
Task 5: Show the top five movies with the highest percentage return on investment. bollywood-movies.csv Failed to (1) clean budget correctly, and (2) calculate missing Revenue values based on India and Worldwide.
Task 6: What were the top three lowest scoring matches? Sort in ascending order and show location, local and visitor team names. euroleague-basketball.csv Failed to select related columns for calculating scores (1), and not knowing how to aggregate scores by Game and Round (2).
@@ 5.3. Study Procedure
The order of the three Phasewise, Stepwise, and Conversational (baseline) systems was fully counterbalanced using a Latin square design across the 18 participants and tasks to minimize order effects, while tasks were fixed from T1 to T6. Participants spent approximately 50 minutes with each system. They received a 10-minute tutorial and a 5-minute warm-up task to familiarize themselves with the system. Then proceeded to the main study tasks, where they were given a dataset and a NL query. The lead author who conducted all experiments, explained the dataset and relevant columns for each task. Participants proceeded to execute the task and were asked to think aloud throughout the study (Fonteyn et al., 1993).
Participants were made aware that identifying and correcting mistakes made by the AI was their responsibility. They were instructed to notify the experimenter once they believed they have achieved a correct result using the AI tool. The experimenter would then verify their solution against expected outcomes and provide up to two hints if necessary. These hints addressed AI mistakes correlating with each of the two issues for each task, as listed in Table 2, ensuring consistency across participants. Completion criteria for each task required resolving both issues listed in the table.
Following the completion of each task, participants were asked about their choice of method for steering the AI (e.g., editing the execution plan versus directly editing the code) and their verification processes. After completing two tasks under each condition, participants completed a questionnaire including Likert items about their ability to verify, intervene and steer the AI, sense of control, information overload, frustration levels, and the utility of specific features (exact questions included in Figure 9). Additionally, participants discussed their experience with each system in a 5-minute semi-structured interview.
The sessions were conducted in-person, lasting approximately 2.5 hours with a short break after using each system. Consent was obtained before running the study and each participant was compensated with a GBP £50 Amazon gift card. Our study protocol was reviewed and approved by our institution’s ethics and compliance review board.
@@ 5.4. Data Collection and Analysis
We recorded the audio and screen activity during each session using MS Teams. Audio recordings were transcribed for analysis. User interactions and feature usage was also logged.
The think-aloud data was our main source of understanding how participants used the different systems and what they thought about them in comparison with each other. We transcribed the think-aloud data and post-condition interviews, and two researchers performed a negotiated, directed qualitative analysis. Ahead of the analysis, we identified a set of research themes concerning steering and verification during the AI-assisted data analysis process, and report our findings organised by these themes in Section 6. Because we were interested in specific themes a priori, and were not developing a reusable coding scheme, our analysis differs from the commonly applied inductive approach (Braun and Clarke, 2006). We did not develop a codebook, and this is not a situation in which it is appropriate to seek inter-rater reliability. Instead, we used a “deductive” coding approach focusing on steering and verification as the main themes. The two researchers iteratively discussed their interpretation of the findings and negotiated each disagreement until it was resolved (McDonald et al., 2019).
@@
We analyzed post-condition Likert responses to compare the three systems and determine any statistically significant differences using a Friedman Chi Square test on the responses for each question with the system type as the independent variable. When significant differences were found ($\alpha=0.05$), a Wilcoxon signed-rank test identified pairwise significant comparisons. Since we made three comparisons (between each pair of systems), we applied a Bonferroni correction ($\alpha=0.016$).
@@ 6. Results
In this section, we present a comparative analysis of the Stepwise and Phasewise AI tools versus the Conversational baseline. Our findings are derived from study observations, log data, participant (P1–P18) think-aloud data, post-condition surveys, and post-study interviews. In turn, we present the results regarding task completion (Section 6.1), steering and control (Section 6.2), and verification (Section 6.3).
@@ 6.2.1. Steering by Directly Editing AI’s Assumptions and Actions
6.2.2. Higher Perceived Control Through Step-by-Step Task Decomposition
A distinct advantage observed with the Stepwise system was the enhanced control participants reported over the data analysis process. This perception was mainly attributed to tackling the task in smaller, manageable segments. For instance, P16 reported an increased sense of control, by being able to “easily edit the assumptions and actions in each step.” Similarly, P11 felt “much happier” and “more confident”, attributing it to the ability to “manipulate steps naturally.” P10 also shared a sense of more control over the AI, stating that they “were not scared” to make edits to what the AI was doing, as “it was just a couple of lines,” and “it was more inviting to edit the assumptions.” P17 mentioned that the Stepwise system facilitated an iterative process “where [they] could easily go back and change something”. While these reports indicate a perception of enhanced control, actual control should be measured in future work.
@@ 6.2.3. Steering by Manually Editing Code
Most participants appreciated the ability to manually edit AI-generated code. P18 mentioned editing the AI-generated code was like “you’re taking over from the AI.” These edits ranged from minor modifications, such as manually changing a threshold or printing values, to more involved changes such as using the Generate code feature to update the logic behind a line of code. When using the Generate code feature, participants (P1, P3, P4, P7, P8, P9 P11, P12, P17, P18) selected a line of code and prompted the AI to update it based on a provided natural language query. For instance, P9 selected df[df[’company_location’] == US’] in their code and prompted the AI with ‘‘can you change this line to look for containing ‘US’ instead of strict equality?’’ P12 experienced increased control when using the Generate code feature since they “could make very granular prompts”.
@@ 6.2.4. Preference for Conversational Steering
In some instances, participants (P4, P5, P6, P8, and P11) specifically indicated a preference for steering the AI through natural language prompts. They favored the Conversational method of steering the AI over editing the structured assumptions, actions, or execution plans. For example, P4 expressed difficulty in understanding how changes to the assumptions in the Phasewise system, affected the LLM’s output. In contrast, P4 had a more accurate mental model of how to interact with the Conversational baseline, stating “I know exactly how writing a prompt is going to affect it.” Others felt constrained by the need to adhere to a specific structure, expressing a preference for more free-form interactions. For instance, P8 described the Conversational system as easier and faster for “directing the AI using natural language”, compared to the Stepwise system where they felt they were “trying to change the syntax of the AI.” Similarly, P5 wanted to “intentionally write vague prompts and see how much [the AI] understands.” P4 believed that the Conversational system required less cognitive effort, stating “I don’t like spending that effort to think about it.” P11 mentioned feeling “less critical” about themselves when using the Conversational tool, allowing the AI to “go and figure it out” on their behalf. P11 elaborated: “I knew what it was [that] I wanted it to consider, but when [the tool] is expecting a structured input then I was more concerned with providing it in a nice and structured manner.”
@@ 6.2.5. Avoiding Edits that Lead to Inconsistency or Regeneration
  • if they had previously edited downstream components and then decided to update an upstream component, the system ignored all downstream edits since regenerations proceed from top to bottom.
  • if they decided to make edits to downstream components that conflicted with assumptions or code in upstream components.
@@ 6.3.1. Verification through Reading Code and Asking Questions
Reading the code line-by-line was a common verification method. P12 mentioned that “you still have to read all the code and understand what it’s doing” for verification. Participants mostly relied on their own knowledge about Python and Pandas for verification, as stated by P8: “you have to know how to code, because you have to read the code and make sure it makes sense.” When P3 was asked how they knew that they had successfully finished the task, they responded “I inspected the code and found that it handled that edge case correctly.” However, in many instances participants had difficulty understanding the AI-generated code, if it used idioms or functions unfamiliar to the participant. Participants appreciated the Ask Question feature in these situations. A majority of participants (n=13) used this feature at least once to explain a portion of the generated code. For example, when P9 was working on Task 2, they stated: “I’ve never seen this function explode() so I’m just gonna ask what does this do”’. Participants found the responses to their queries useful. For example, P2 asked about the fillna() function and realized that the code was doing something undesirable (replacing nan with "Unknown"). Furthermore, participants felt the absence of the Ask Question feature when using the baseline tool, where for instance P18 wanted to use a search engine, and P6 mixed the main thread of the task with a comprehension question: “I don’t understand [refers to code].” During the study, several unanticipated, yet effective use cases of the Ask Question feature emerged, such as P4 asking questions about an assumption, and P12 requesting help with debugging by selecting part of the code and asking “why this code produces an error.”
@@ 6.3.2. Inspecting Intermediary Variables
All participants used the intermediary variable inspection feature available in all three systems for verification. Participants inspected variables to “compare between turns” (P1), and to “see if [the system] has done the [operation] correctly” (P14). P12 mentioned that the verification process was similar to “debugging” and P11 indicated that inspecting all the dataframes significantly increases confidence in the process.